OpenStudy (anonymous):

find the missing coordinate to complete the ordered pair (3, ) so that satisfies the equation:3x-y=4

jimthompson5910 (jim_thompson5910):

(3, ) means x = 3 and y is unknown

jimthompson5910 (jim_thompson5910):

Plug in x = 3 and solve for y 3x-y=4 3*3-y=4 9-y = 4 9 = 4+y 9-4 = y 5 = y y = 5 So the ordered pair, ie point, that is on this line is (3,5)